OpenStudy (anonymous):

A heavy crate applied a force of 2,000 N on a 50-m2 piston. What force needs to be exerted on a 1.00 m2 piston to lift the crate?

OpenStudy (amistre64):

i beleive: force times surface area should equal or is it force/sa ?

OpenStudy (amistre64):

at any rate, there is a ration of Fa/Fb equal to the ratio of the surface areas

OpenStudy (anonymous):

So How would I go about solving it? And strangely (or sadly) enough I wasn't given a formula!

OpenStudy (amistre64):

its similar in thinking to a seesaw; the fat kids have to sit closer to balance out the skinny kid at the end

OpenStudy (amistre64):

f1/f2 = sa1/sa2

OpenStudy (amistre64):

you know f1 sa1 and sa2

OpenStudy (anonymous):

I think I do but I forgot what fl sal and sa2 stand for

OpenStudy (amistre64):

f for force, sa for surface area ...

OpenStudy (anonymous):

OK So to solve for the question A heavy crate applied a force of 2,000 N on a 50-m2 piston. What force needs to be exerted on a 1.00 m2 piston to lift the crate?

OpenStudy (anonymous):

So do I divide 1.00 m2 by 50-m2?

OpenStudy (anonymous):

errr times

OpenStudy (amistre64):

\[\frac{F_1}{SA_1}=\frac{F_2}{SA_2}\] \[F_1\frac{SA_2}{SA_1}=F_2\]

OpenStudy (amistre64):

2000*1/50

OpenStudy (anonymous):

OH OK!

OpenStudy (anonymous):

One second

OpenStudy (anonymous):

A heavy crate applied a force of 2,000 N on a 50-m2 piston. What force needs to be exerted on a 1.00 m2 piston to lift the crate? So.. 2,000N* 1m2/50m2=40

OpenStudy (anonymous):

Right?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

What is the unit of 40?

OpenStudy (amistre64):

what is: N m^2/m^2 ??

OpenStudy (anonymous):

I wish I knew....

OpenStudy (anonymous):

Unit's confuse me

OpenStudy (anonymous):

I think it's Nm^2

OpenStudy (amistre64):

think of them as variables instead

OpenStudy (amistre64):

no, if the Force on the left is in Newtons, then the force on the right will be in Newtons ....

OpenStudy (anonymous):

So That means it is N?

OpenStudy (amistre64):

but regardless of the fact .. N m^2/m^2 , anything divided by itself is equal to 1 N*1 = N

OpenStudy (anonymous):

Because the m^2 cancel each other out

OpenStudy (anonymous):

So my final answer should be 40N?

OpenStudy (amistre64):

thats what all the parts togheter form ... yes. assuming i recall this material correctly :)

OpenStudy (anonymous):

Thanks @amistre64 :D

OpenStudy (amistre64):

youre welcome